# High-dimensional dynamics in low-dimensional networks

### v4

@@ Results
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+{\bf{F}}({\bf{z}},\bm{\eta})$ (4)
for some smooth vector field, ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$, where $\tau>0$ sets the timescale of the dynamics. Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and $\bm{\eta}(t)$ is an $N\times 1$ vector of external perturbations, inputs, and/or noise.
In general, the dependence of ${\bf{F}}({\bf{z}},\bm{\eta})$ on $\bf{z}$} defines interactions between nodes, but it is not immediately clear how to define a fixed network of interactions, $W$, from the dynamics in Eq. (4). When dynamics are linear, $\bf{F}$} can be written as ${\bf{F}}({\bf{z}},\bm{\eta})=W{\bf{z}}+{\bf{G}}(\bm{\eta})$, and we can take $W$ to represent the effective connectivity. In nonlinear systems, the effective connectivity between nodes can change with the state of the system. To resolve this difficulty in defining network structure when dynamics are nonlinear, we assume that dynamics represent a small deviation away from a stable equilibrium so that we can linearize them.
Specifically, we assume that there exists a time-constant forcing, $\bm{\eta}(t)=\bm{\eta}_{0}$, that produces a stable steady-state solution, ${\bf{z}}_{0}$. In other words, there exist $\bm{\eta}_{0}$ and ${\bf{z}}_{0}$ such that
${\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})={\bf{z}}_{0}.$
$J=-I+\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$
has eigenvalues with strictly negative real part. Here, $I$ is the identity matrix and $\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ is the Jacobian matrix of $\bf{F}$} with respect to $\bf{z}$} evaluated at the fixed point. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
$\bm{\eta}_{p}(t)=\bm{\eta}_{0}+\epsilon\bm{\eta}(t).$
The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, $\bm{\eta}_{p}(t)$, can be written to linear order in $\epsilon$ as
${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$ (5)
The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+{\bf{x}}(t)$ (6)
where
$W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ (7)
is interpreted as the effective connectivity matrix and
${\bf{x}}(t)=\partial_{\bm{\eta}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})\bm{\eta}(t)$
is interpreted as the effective input. Our assumption that the equilibrium is stable implies that all eigenvalues of $W$ have real part less than $1$.
In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\bm{\eta})={\bf{H}}({\bf{z}})+\bm{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\bm{\eta}$. In other models, the raw input, $\bm{\eta}$, to the nonlinear system in Eq. (4) can have a different structure and dimensionality to the effective input, $\bf{x}$}, to the linearized system in Eq. (6), so care must be taken in interpreting our results. For illustrative purposes, we consider a contrived example of such a model in Figure 11 in Appendix D.
Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix. For example, the system might be written as $d{\bf{z}}/dt={\bf{H}}(W{\bf{z}})+{\bf{x}}$ where $W$ is meant to quantify a connectivity matrix, but $\partial_{\bf{z}}{\bf{H}}(W{\bf{z}}_{0})\neq W$. In this case, our interpretation of the effective connectivity matrix is not consistent with the original parameterization, but they will often share structural properties such as their approximate dimensionality. We consider one such example from epidemiology at the end of the Results.
The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive. However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Figures 13 and 14 in Appendix D for two examples of nonlinear, unstable dynamics). Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state. Previous work [1] considered models of the form $d{\bf{z}}/dt={\bf{G}}({\bf{y}},{\bf{z}})$ where ${\bf{y}}=W{\bf{z}}$ (note that this model lacks external input). This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$. We discuss the relationships between their analysis and ours in more depth in Appendix A.
In summary, we assume that the network dynamics are either fully linear or, if they are nonlinear, they represent a small perturbation away from a stable equilibrium. This assumption allows us to focus our analysis on the linearized dynamics in Eq. (6).
Assumption 2: The connectivity matrix is strongly low rank in an asymptotic sense.
We now formalize an asymptotic notion of a strongly low-rank network. Specifically, we consider the limit of large network size, $N\to\infty$, and assume that $W$ has $r\sim\mathcal{O}(1)$ asymptotically large singular values,
$\sigma_{k}\gg 1,\;\;k=1,\ldots,r$
where $\sigma_{k}$ denotes the $k$th largest singular value of $W$. We assume that the remaining singular values are small or moderate in magnitude,
$\sigma_{j}\leq\mathcal{O}(1),\;\;j=r+1,\ldots,N.$
Later, we consider some examples in which these scaling laws arise naturally.
Matrices that satisfy the assumptions above can be decomposed as [1]
@@
where $W_{0}$ has rank $r$ and asymptotically large singular values, while the singular values of $W_{1}$ are $\mathcal{O}(1)$ at most and $W_{1}$ can be full rank. Since $W_{0}$ has rank $r$, we can write its singular value decomposition as
@@
where $\Sigma$ is a diagonal, $r\times r$ matrix of singular values ($\Sigma_{kk}=\sigma_{k}\gg 1$) while $U=[{\bf{u}}_{1}\cdots{\bf{u}}_{r}]$ and $V=[{\bf{v}}_{1}\cdots{\bf{v}}_{r}]$ are $N\times r$ orthonormal matrices with columns ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$ that define the left- and right-singular vectors of $W_{0}$.
We additionally assume that $W_{1}$ is a random matrix that is independent of $U$ and $V$. In other words, $W$ is a low-rank matrix perturbed by noise. This is a stronger assumption than necessary ($W_{1}$ need not be fully independent from $W_{0}$ for our results to hold) but it simplifies our analysis.
Strongly low rank structure is a special case of low-rank structure. For example, a network could have a small number of $\mathcal{O}(1)$ singular values with the remaining singular values $o(1)$. Such a network would still be effectively low-rank in an asymptotic sense, but it would not be strongly low-rank according to our definition because the dominant singular values are not asymptotically large. These “weakly low-rank” networks [2] can exhibit a weaker form of low-rank suppression and high-dimensional dynamics, depending on the details of their structure (see Appendix A for further discussion and examples).
Low-rank suppression occurs under Assumptions 1 and 2.
Our first result is that Assumptions 1 and 2 are sufficient for low-rank suppression (as observed in Figures 1d and 2d). Specifically,
Claim 1.
Consider the dynamics defined by Eq. (6) under Assumptions 1 and 2. There exists an input ${\bf{x}}_{a}$ with $\|{\bf{x}}_{a}\|=\mathcal{O}(1)$ such that the network’s steady-state response to ${\bf{x}}_{a}$ is asymptotically larger than the $\mathcal{O}(1)$ response to a random input, ${\bf{x}}_{\textrm{r}}$ (where $\|{\bf{x}}_{\textrm{r}}\|=\mathcal{O}(1)$).
To show why Claim 1 is true, first choose a $k\in\{1,\ldots,r\}$ and define
@@
Note that $\|{\bf{x}}_{a}\|=1+o(1)$ since $\sigma_{k}\gg 1$, $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, and the maximum singular value of $W_{1}$ is $\mathcal{O}(1)$. Now note that the steady-state response to the input ${\bf{x}}_{a}$ is given by
@@
This can be checked by substituting the equations for ${\bf{x}}_{a}$ and ${\bf{z}}_{a}$ into the steady-state equation, ${\bf{z}}=W{\bf{z}}+{\bf{x}}$. Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$. Since $\sigma_{k}\gg 1$, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq. (8). In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is normal. This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
@@
Note from Eq. (8) that ${\bf{x}}_{a}$ is closely aligned to ${\bf{u}}_{k}$ since $\sigma_{k}\gg 1$. Hence, the suppressed input, ${\bf{x}}_{a}$, is nearly (but not exactly) parallel to the left singular vector, ${\bf{u}}_{k}$. Also note that the response, ${\bf{z}}_{a}$, is perfectly aligned to ${\bf{v}}_{k}$. Finally, note that there are $r$ different choices for ${\bf{x}}_{a}$ (one for each $k\in\{1,\ldots,r\}$), all of which are nearly orthogonal to each other.
@@
Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Figure 10 in Appendix D for which ${\bf{z}}=W{\bf{x}}$ in the steady-state). More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$). This is a left-right reversal and a suppression-amplification reversal of the behavior for recurrent networks, for which inputs aligned to ${\bf{u}}_{k}$ produce suppressed responses in the direction of ${\bf{v}}_{k}$.
@@
In summary, recurrent networks with strongly low-rank structure satisfying Assumptions 1 and 2 admit a small number of input directions along which inputs are strongly suppressed. We call this form of suppression low-rank suppression. In later sections, we will explore consequences of low-rank suppression in common low-rank network structures. Before discussing the dimensionality of dynamics, we first need to make an additional assumption about the dimensionality of the input.
@@
Assumption 3: External inputs are high-dimensional and slowly varying.
To begin discussing the dimensionality of ${\bf{z}}(t)$, we need a model of ${\bf{x}}(t)$ that allows us to quantify dimensionality in a concrete way. To this end, we assume that ${\bf{x}}(t)$ is a stationary ergodic stochastic process in which case ${\bf{z}}(t)$, as defined in Eq. (6), is also a stationary ergodic process whenever dynamics are stable [31]. This allows quantify the dimensionality of ${\bf{x}}(t)$ and ${\bf{z}}(t)$ in terms of their principal components or, equivalently, by the eigenvalues of their stationary covariance matrices.
We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Appendix B for more mathematical details).
@@
The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics. When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Figure 11 in Appendix D for an example).
@@
We additionally assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions. To make this assumption more precise, note that whenever ${\bf{x}}(t)={\bf{x}}$ is constant, solutions, ${\bf{z}}(t)$, converge to the steady-state given by ${\bf{z}}=[I-W]^{-1}{\bf{x}}$. We assume that ${\bf{x}}(t)$ varies sufficiently slowly that ${\bf{z}}(t)$ is approximated by the quasi-steady state approximation,
${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$
In Appendix B, we give a more mathematically detailed definition of this assumption. In Appendix C, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
In simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d. Gaussian process with correlation time $\tau_{x}=10$. Our assumption of high-dimensionality is satisfied because each ${\bf{x}}_{j}(t)$ is independent. Our assumption that ${\bf{x}}(t)$ varies slowly is satisfied because $\tau_{x}$ is much larger than the network’s intrinsic timescale, $\tau=1$.
Conditions for low-dimensional dynamics.
In Appendix B, we prove that (under Assumptions 1–3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
$P=V^{T}U$

### v5

@@ IV General conditions for low-rank suppression and high-dimensional dynamics
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+{\bf{F}}({\bf{z}},{\bf{x}})$ (4)
for some smooth vector field, ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$, where $\tau>0$ sets the timescale of the dynamics. The dependence of ${\bf{F}}({\bf{z}},\boldsymbol{x})$ on $\bf{z}$} defines interactions between nodes, but it is difficult to define a fixed network of interactions, $W$, from the general dynamics in Eq. (4). In nonlinear systems, the effective connectivity between nodes can change with the state of the system. Moreover, nonlinear recurrent networks can produce any dynamics (high- or low-dimensional) in response to any input [37, 38], so it is difficult to make any general statements about the dimensionality and structure of dynamics in recurrent networks far from a stable equilibrium. To resolve this difficulty, we assume that dynamics represent a small deviation away from a stable equilibrium so that we can linearize them. A change of coordinates under this linearization gives Eq. (1) where $W$ is defined as the Jacobian of $\bf{F}$} at the fixed point ($W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$; see Appendix B for details). Our assumption of stability guarantees that all eigenvalues of $W$ have real part less than 1. To avoid systems close to instability, we additionally assume that the eigenvalues are bounded away from 1, in other words $W$ does not have any eigenvalues close to $1$. For our examples, this implies that the spectral radius, $\rho$, of the random part, $W_{1}$, is not close to $1$. Eigenvalues near $1$ or eigenvalues larger than 11 can produce low-dimensional dynamics through a different mechanism than the mechanisms that we are studying [22, 21, 2, 23, 1] (see Figures 1012 in Appendix F for examples).
IV.2 Assumption 2: The connectivity matrix is strongly low rank in an asymptotic sense.
We additionally assume that the network, $W$, has a strongly low-rank structure. To make this assumption precise, we consider networks in the limit of large $N$ and assume that a small number of singular values diverge with $N$ while the rest remain $\mathcal{O}(1)$ or smaller (see Appendix C for a more precise statement).
This assumption has two components. First, that the connectivity matrix has a small number of dominant singular values. In later sections, we show that this property arises naturally in several natural network structures such as networks with biased weights, modular structure, or smooth spatial structure.
Secondly, this assumption implies that the dominant singular values grow large with $N$. This contrasts to other work on “weakly” low-rank networks in which the largest singular values are $\mathcal{O}(1)$ [3, 7, 5, 8, 9, 10]. Weakly low-rank structure arises naturally when mean weights scale like $\mathcal{O}(1/N)$ whereas strongly low-rank structure arises when mean weights are asymptotically larger than $1/N$, for example when they scale like $\mathcal{O}(1/\sqrt{N})$ as in balanced networks [29, 39, 40, 41, 4]. See Appendix A for a more in-depth comparison between weakly and strongly low-rank networks.
Strongly low-rank networks can be decomposed as [1]
@@
where $W_{0}$ has rank $r\sim\mathcal{O}(1)$ and large singular values, while the singular values of $W_{1}$ are $\mathcal{O}(1)$ at most and $W_{1}$ can be full rank. We specifically assume that $W_{1}$ is a random matrix and independent from $W_{0}$. Since $W_{0}$ has rank $r$, we can write its singular value decomposition as
@@
where $\Sigma$ is a diagonal, $r\times r$ matrix of singular values
$\sigma_{k}=\Sigma_{kk}\gg 1$
while $U=[{\bf{u}}_{1}\cdots{\bf{u}}_{r}]$ and $V=[{\bf{v}}_{1}\cdots{\bf{v}}_{r}]$ are $N\times r$ orthonormal matrices with columns ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$ that define the left- and right-singular vectors of $W_{0}$.
IV.3 Low-rank suppression occurs under Assumptions 1 and 2.
Our first result is that Assumptions 1 and 2 are sufficient for low-rank suppression (as observed in Figures 1d and 2d). Specifically, we claim that low-rank suppression is realized under a static input of the form
@@
where $k\leq r$ so that $\sigma_{k}\gg 1$ is one of the dominant singular values. We claim that inputs of this form are strongly suppressed by the network. To see why this is true, first note that $\|{\bf{x}}_{a}\|=1+o(1)$ since $\sigma_{k}\gg 1$, $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, and the maximum singular value of $W_{1}$ is $\mathcal{O}(1)$. Now note that the steady-state response to the input ${\bf{x}}_{a}$ is given by
@@
This can be checked by substituting the equations for ${\bf{x}}_{a}$ and ${\bf{z}}_{a}$ into the steady-state equation, ${\bf{z}}=W{\bf{z}}+{\bf{x}}$. Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$. Therefore, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq. (5) in the sense that $\|{\bf{z}}_{a}\|\ll\|{\bf{x}}_{a}\|$.
@@
Note from Eq. (5) that ${\bf{x}}_{a}$ is closely aligned to ${\bf{u}}_{k}$ since $\sigma_{k}\gg 1$. Hence, the suppressed input, ${\bf{x}}_{a}$, is nearly (but not exactly) parallel to the left singular vector, ${\bf{u}}_{k}$. Also note that the response, ${\bf{z}}_{a}$, is perfectly aligned to ${\bf{v}}_{k}$. Hence, strongly low-rank recurrent networks suppress inputs aligned near the dominant left singular vectors (${\bf{u}}_{k}$) and the suppressed response is aligned with the dominant right singular vectors (${\bf{v}}_{k}$).
@@
Interestingly, these input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Figure 7 in Appendix F for which ${\bf{z}}=W{\bf{x}}$ in the steady-state). More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$). This is a left-right reversal and a suppression-amplification reversal of the behavior for recurrent networks, for which inputs aligned to ${\bf{u}}_{k}$ produce suppressed responses in the direction of ${\bf{v}}_{k}$.
@@
In summary, recurrent networks with strongly low-rank structure satisfying Assumptions 1 and 2 admit a small number of input directions along which inputs are strongly suppressed. We call this form of suppression low-rank suppression. Low rank suppression occurs regardless of whether connectivity is normal, explaining why low-rank suppression was observed in both Figure 1 and Figure 2. Later, we will explore consequences of low-rank suppression in common network structures. First, we discuss conditions for the separate but related phenomenon of high-dimensional dynamics.
@@
IV.4 Assumption 3: External inputs are stationary, high-dimensional, and slowly varying.
To begin discussing the dimensionality of ${\bf{z}}(t)$, we need to make assumptions that allow us to define precisely what we mean by high- and low-dimensional dynamics. To this end, we assume that ${\bf{x}}(t)$ is a stationary ergodic stochastic process. In this case ${\bf{z}}(t)$ is also a stationary ergodic process whenever dynamics are stable [42] as in Assumption 1. This allows us to quantify the dimensionality of ${\bf{z}}(t)$ in terms of its principal components or, equivalently, by the singular values of its stationary covariance matrix. Specifically, the variance explained by the first $k$ principal components represents the amount of variability constrained to a $k$-dimensional linear manifold. For this reason, we define ${\bf{z}}(t)$ to have low-dimensional dynamics when there exists an $r\sim\mathcal{O}(1)$ such that $\sigma_{j}/\sigma_{k}\to 0$ as $N\to\infty$ whenever $k\leq r$ and $j>r$. This implies that the vast majority of variability lies on a low-dimensional linear manifold. Otherwise, we say that ${\bf{z}}(t)$ has high-dimensional dynamics.
@@
We also assume that ${\bf{x}}(t)$ is high-dimensional. If ${\bf{x}}(t)$ were low-dimensional, then low-dimensional dynamics in ${\bf{z}}(t)$ could be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Figure 8 in Appendix F for an example). In assuming that ${\bf{x}}(t)$ is high-dimensional, we are focusing on the question of whether low-dimensional dynamics arise from network interactions, not from low-dimensional input alone.
For simplicity, within this section, we specifically assume that the components, ${\bf{x}}_{j}(t)$, are i.i.d. This implies that ${\bf{x}}(t)$ is high-dimensional, is a stronger assumption than necessary, but simplifies the exposition. In Appendix D, we consider more general inputs that are not assumed to be i.i.d.
@@
For now, we also assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions. To make this assumption more precise, note that whenever ${\bf{x}}(t)={\bf{x}}$ is constant, solutions, ${\bf{z}}(t)$, converge to the steady-state given by ${\bf{z}}=[I-W]^{-1}{\bf{x}}$. We assume that ${\bf{x}}(t)$ varies sufficiently slowly that ${\bf{z}}(t)$ is approximated by the quasi-steady state approximation,
${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$ (6)
In Appendix D, we give more details about this assumption. In Appendix E, we extend our analysis and results to the case that ${\bf{x}}(t)$ is not slowly varying.
IV.5 Conditions for low-dimensional dynamics under Assumptions 1–3
Under the quasi-steady state approximation in Eq. (6) and the assumption that elements of ${\bf{x}}(t)$ are i.i.d., it is easy to show that the covariance matrix of ${\bf{z}}(t)$ can be written as [42]
$\overline{C}^{\bf{z}}\propto[I-W]^{-1}[I-W]^{-T}.$
As a result, the question of whether ${\bf{z}}(t)$ is low-dimensional becomes a question of whether $I-W]^{-1}$ has a small number of dominant singular values whenever $W$ has a small number of dominant singular values. More details are given in Appendix D where we prove that the answer to this question depends on the recurrent alignment matrix [4] of $W_{0}$, defined by
$P=V^{T}U.$